OpenStudy (anonymous):

Which description best describes the solution to the following system of equations? y = –x + 4 y = 3x + 3

OpenStudy (anonymous):

Line y = –x + 4 intersects the line y = 3x + 3. Lines y = –x + 4 and y = 3x + 3 intersect the x-axis. Lines y = –x + 4 and y = 3x + 3 intersect the y-axis. Line y = –x + 4 intersects the origin.